# Particle Multi-Axis Transformer for Jet Tagging

### v1

@@ Abstract
Jet tagging is an essential categorization problem in high energy physics. In recent times, Deep Learning has not only risen to the challenge of jet tagging but also significantly improved its performance. In this article, we propose an idea of a new architecture, Particle Multi-Axis transformer (ParMAT) which is a modified version of Particle transformer (ParT). ParMAT contains local and global spatial interactions within a single unit which improves its ability to handle various input lengths. We trained our model on JETCLASS, a publicly available large dataset that contains 100M jets of 10 different classes of particles. By integrating a parallel attention mechanism and pairwise interactions of particles in the attention mechanism,ParMAT achieves robustness and higher accuracy over the ParT and ParticleNet. The scalability of the model to huge datasets and its ability to automatically extract essential features demonstrate its potential for enhancing jet tagging.
@@ 1 Introduction
2 Related Work
Deep Learning for Jet tagging: Deep learning has become the go-to approach for jet tagging as it has significantly improved particle identification in recent years. The efficiency and effectiveness of DL techniques on jet physics has improved with the advancement of two factors.
@@ 2 Related Work
Deep-learning-powered jet tagging techniques have become integral to data analysis at the LHC, notably through innovations like the DeepAK8 algorithm developed by the CMS Collaboration[15]. This algorithm, employing a 1D CNN[16] based on the ResNet architecture[17], effectively identifies jets arising from top quarks or bosons like the Higgs, W, or Z. Such advancements have markedly heightened the potential for discovering new heavy particles, as evidenced by CMS publications in 2021[18] and 2022[19]. Additionally, ParticleNet, another deep-learning tool utilized by CMS, enabled groundbreaking achievements, including the first observation of Z boson decay into charm quarks and setting stringent constraints on Higgs decay to charm quarks[19]. Furthermore, ParticleNet’s application in probing the quartic Higgs and vector bosons interaction has led to the indirect confirmation of its existence. These strides underscore the crucial role of improved jet tagging methods in advancing our knowledge of elementary particles, the fundamental constituents of the universe.
@@ 3 Dataset
4 Model Architecture
@@
This paper introduces a modified transformer-based architecture called Particle Multi-axis Transformer(ParMAT), a combination of Particle Transformer and Multi-axis Transformer. We split attention into two types: local and global.[3] Our approach uses layers of block and grid attention in sequence. Block region operates within discrete, non-overlapping windows (small patches on intermediate feature maps) to identify local patterns. In contrast, grid region targets long-range (global) interactions by focusing on a uniformly spaced, sparse grid. The sizes of the windows for both block and grid attentions are adjustable parameters, allowing us to maintain linear computational complexity relative to the size of the input. It serves as a core component facilitating both local and global spatial interactions within a single unit.ParMAT offers increased adaptability and effectiveness compared to full self-attention, showcasing its ability to handle various input lengths with linear complexity. Unlike traditional local attention mechanisms, ParMAT expands model capacity by incorporating a global receptive field. ParMAT is constructed by stacking a particle attention multi-axis block, particle multi-head attention,multi-axis block attention,multi-axis grid attention and class attention mechanism. ParMAT is a modified version of ParT inspired by attention mechanism.[37]. The model of this architecture is shown in Figure 2(a). We consider $N$ number of particles in a jet.
MLP Projection: Each set of inputs is passed through a Multi-Layer Perceptron (MLP) to project the $R$ and $R^{\prime}$ features into $n$ and $n^{\prime}$-dimensional embeddings, $x_{0}\in\textbf{R}^{N\times n}$ and $W\in\textbf{R}^{N\times N\times n^{\prime}}$ for $particles$ and $Interactions$ respectively, where $x0$ and $W$ are the inputs after an MLP has processed them. As the order in which particles in a jet are processed does not matter we do not need any positional encoding. So we feed $x_{0}$ into multiple Particle Multi-Axis Attention blocks. Each block output is fed into the next block as an input. This produces a series of new embeddings [38].
@@
Particle Multi-Axis attention block: In Figure 2(b), the P-MAA block consists of two parts. First part consists of a LayerNorm (LN) then a multi-head attention (MHA) module and two Multi-axis attention, MAA (grid/global) and MAA (Window/local) with a LayerNorm (LN) layer again. The the second part includes the 2-layers of Multi-layer Perceptron having the LayerNorm (LN) before it and GELU nonlinearity in between. Each layer is connected with the following one. The overall block structure is taken from NormFormer. we modified MHA as P-MHA (particle-Multi head attention), which deals with the pairwise particle interaction. The P-MHA is computed as
@@
where, Q, K, and V are the query, key, and value. These are the linear projections of the particle embedding in our case $x^{L}$. We modify the pre-softmax attention weights by adding the interaction matrix W. This increases the expressiveness of the attention mechanism by enabling PMHA to alter the dot-product attention weights and integrate particle interaction information.
@@
Class Attention Block: The class attention block as illustrated in figure 2(b) has almost similar architecture as the particle multi-axis attention block. In this block, we use simple MHA and calculate attention between particles and Global class token $x_{class}$. MHA inputs are:
@@
Model Implementation: We implement our ParMAT model in PyTorch[40]. Particularly, P-MHA and P-MAA is implemented using the PyTorch. The ParMAT model includes 8 P-MAA blocks and 2 class attention blocks. we provide interaction matrix $W$ as the attentionn mask input. It utilizes a 128-dimensional particle embedding, generated from the input particle features through a 3-layer MLP with (128, 512, 128) nodes in each layer, utilizing GELU nonlinearity, and LN for normalization. The interaction input features are processed through a 4-layer pointwise 1D convolution with channel configurations (64, 64, 64, 8), GELU nonlinearity, and batch normalization to create an 8-dimensional interaction matrix. The P-MHA (MHA) in the particle (class) attention blocks have 8 heads, each with a query dimension of 16 and an expansion factor of 4 for the MLP. We use a dropout (a regularization technique used in neural networks to prevent overfitting) of 0.1 for all P-MAA blocks and no dropout for the class attention block.
5 Experiment
We perform experiments using the new JETCLASS dataset, a widely adopted benchmark for jet tagging tasks in high-energy physics research. The dataset comprises simulated particle collision events, with each event containing detailed information about the particles produced in the collision, including their momentum and energy. The results of our experiment are summarized in Section VIII, where we also compare the performance to previous leading methods.
@@ 6 Evaluation Metrices
In the context of jet tagging, separating signal jets (class s) from background jets (class b) is crucial. This equation, inspired by the methodologies of the CMS experiment at the LHC, utilizes softmax outputs (Score(s) and Score(b)) to achieve optimal discrimination between the two classes. Background rejection is challenging metric in jet tagging. This is due to significant improvement in background rejection directly translates to a substantial increase in the LHC’s discovery potential. For example, doubling the background rejection can result in a 40% increase in discovery potential, which is comparable to either doubling the dataset size or operating the LHC for twice the duration.
 [^3] The ROC AUC function available in scikit-learn library may be used to calculate the area under the curve by passing in the arguments, average equals to ’macro’ and multi-class equals to ’ovo’.
7 Results
@@
Evaluation metrics described in Section VII are used to determine the performance of our proposed model on JETCLASS dataset, with the outcomes compiled in Table 1.This new proposed architecture sets a new benchmark by achieving the highest performance of 86.32%, surpassing all others. ParMAT performs with higher accuracy than Particle transformer.ParMAT significantly outperform the earlier ParT,ParticleNet and P-CNN models on this comprehensive dataset, showing background rejection differences that may exceed an order of magnitude. This significant enhancement in the Particle Multi-Axis Transformer is projected to significantly enhance the discovery capabilities for physics research at the LHC.
@@
Figure 3: Training Accuracy vs Epochs.
Figure 4: Validation Accuracy vs Epochs.
@@
8 Conclusion
We proposed Multi-axis particle transformer in this paper for jet tagging task on particle collision data from CMS detector. Our model performed well on JETCLASS dataset by achieving a reasonable accuracy for classifying our signals into 10 classes and surpassed many previous ML models proposed for jet tagging including the four baselines models. The AUC of 98.76% reflects the strong discriminatory power of our model and its robustness. It also shows the model’s ability to correctly identify our signals while minimizing the false positives. It also suggest the ability of our model for jet tagging task where precise classification is crucial and false positives can lead to significant drawbacks. Ranger optimizer performed better than Adam for training this model (Complete details about training parameters have been covered in section VI: Experiment). Background rejection matrix shows that our model significantly improved on background noise rejection for both decay modes of top quarks $(t\xrightarrow{}bqq^{\prime}$ $t\xrightarrow{}blv)$ and one decay mode of Higgs$(H\xrightarrow{}gg)$. This improvement resulted from integration of multi axis attention block which uses multi axis attention mechanism to explore relations between different particles and jets by considering multiple axis at a time. Results support the fact that transformers can indeed be a good solution for classification tasks in particle physics where we have large amounts of data generated by detectors. Furthermore, larger datasets would be more effective for training such models in future[44].
@@ Acknowledgements
We thank Pattern Recognition Lab (PR-Lab) and the Pakistan Institute of Engineering and Applied Sciences (PIEAS) for providing the necessary computational resources and a healthy research environment.
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@ Abstract
Jet tagging is an essential categorization problem in high energy physics. In recent times, Deep Learning has not only risen to the challenge of jet tagging but also significantly improved its performance. In this article, we proposed an idea of a new architecture, Particle Multi-Axis transformer (ParMAT) which is a modified version of Particle transformer (ParT). ParMAT contains local and global spatial interactions within a single unit which improves its ability to handle various input lengths. We trained our model on JETCLASS, a publicly available large dataset that contains 100M jets of 10 different classes of particles. By integrating a parallel attention mechanism and pairwise interactions of particles in the attention mechanism, ParMAT achieves robustness and higher accuracy over the ParT and ParticleNet. The scalability of the model to huge datasets and its ability to automatically extract essential features demonstrate its potential for enhancing jet tagging.
@@ 1.   Introduction
2.   Related Work
Deep Learning for Jet tagging: Deep learning techniques like CNN and Transformers have shown promising results in different domains[4], and these techniques are also gaining popularity for jet tagging as it has significantly improved particle identification in recent years[5][6][7][8]. The efficiency and effectiveness of DL techniques on jet physics has improved with the advancement of two factors.
@@ 2.   Related Work
Deep learning powered jet tagging techniques have become integral to data analysis at the LHC, notably through innovations like the DeepAK8 algorithm developed by the CMS Collaboration[7]. This algorithm, employing a 1D CNN[19] based on the ResNet architecture[20], effectively identifies jets arising from top quarks or bosons like the Higgs, W, or Z. Such advancements have markedly heightened the potential for discovering new heavy particles, as evidenced by CMS publications in 2021[21] and 2022[5]. Additionally, ParticleNet, another deep-learning tool utilized by CMS, enabled groundbreaking achievements, including the first observation of Z boson decay into charm quarks and setting stringent constraints on Higgs decay to charm quarks[5]. Furthermore, ParticleNet’s application in probing the quartic Higgs and vector bosons interaction has led to the indirect confirmation of its existence. These strides underscore the crucial role of improved jet tagging methods in advancing our knowledge of elementary particles, the fundamental constituents of the universe.
@@ 3.   Dataset
4.   Model Architecture
This paper introduces a modified transformer-based architecture called Particle Multi-axis Transformer(ParMAT), a combination of Particle Transformer and Multi-axis Transformer. We split attention into two types: local and global[3]. Our approach uses layers of block and grid attention in sequence. Block region operates within discrete, non-overlapping windows (small patches on intermediate feature maps) to identify local patterns. In contrast, grid region targets long-range (global) interactions by focusing on a uniformly spaced, sparse grid. The sizes of the windows for both block and grid attentions are adjustable parameters, allowing us to maintain linear computational complexity relative to the size of the input. It serves as a core component facilitating both local and global spatial interactions within a single unit. ParMAT offers increased adaptability and effectiveness compared to full self-attention, showcasing its ability to handle various input lengths with linear complexity. Unlike traditional local attention mechanisms, ParMAT expands model capacity by incorporating a global receptive field. ParMAT is constructed by stacking a particle attention multi-axis block, particle multi-head attention, multi-axis block attention, multi-axis grid attention and class attention mechanism. ParMAT is a modified version of ParT inspired by attention mechanism[39]. The model of this architecture is shown in Figure 2(a). We consider $N$ number of particles in a jet.
MLP Projection: Each set of inputs is passed through a Multi-Layer Perceptron (MLP) to project the $R$ and $R^{\prime}$ features into $n$ and $n^{\prime}$-dimensional embeddings, $x_{0}\in\textbf{R}^{N\times n}$ and $W\in\textbf{R}^{N\times N\times n^{\prime}}$ for particles and Interactions respectively, where $x0$ and $W$ are the inputs after an MLP has processed them. As the order in which particles in a jet are processed does not matter, we do not need any positional encoding. So we feed $x_{0}$ into multiple Particle Multi-Axis Attention blocks. Each block output is fed into the next block as an input. This produces a series of new embeddings[40].
@@ 4.   Model Architecture
Particle Multi-Axis attention block: In Figure 2(b), the P-MAA block consists of two parts. First part consists of a LayerNorm (LN) then a multi-head attention (MHA) module and two Multi-axis attention, MAA (grid/global) and MAA (Window/local) with a LayerNorm (LN) layer again. The the second part includes the 2-layers of Multi-layer Perceptron having the LayerNorm (LN) before it and GELU non-linearity in between. Each layer is connected with the following one. The overall block structure is taken from NormFormer. We modified MHA as P-MHA (particle-Multi head attention), which deals with the pairwise particle interaction. The P-MHA is computed as
@@
where, Q, K, and V are the query, key, and value. These are the linear projections of the particle embedding in our case $x^{L}$. We modify the pre-softmax attention weights by adding the interaction matrix W. This increases the expressiveness of the attention mechanism by enabling P-MHA to alter the dot-product attention weights and integrate particle interaction information.
@@
Class Attention Block: The class attention block as illustrated in figure 2(b) has almost similar architecture as the particle multi-axis attention block. In this block, we used simple MHA and calculate attention between particles and Global class token $x_{class}$. MHA inputs are:
@@
Model Implementation: We implement our ParMAT model in PyTorch[42]. Particularly, P-MHA and P-MAA is implemented using the PyTorch. The ParMAT model includes 8 P-MAA blocks and 2 class attention blocks. We provide interaction matrix $W$ as the attention mask input. It utilizes a 128-dimensional particle embedding, generated from the input particle features through a 3-layer MLP with (128, 512, 128) nodes in each layer, utilizing GELU non-linearity, and LN for normalization. The interaction input features are processed through a 4-layer point wise 1D convolution with channel configurations (64, 64, 64, 8), GELU non-linearity, and batch normalization to create an 8-dimensional interaction matrix. The P-MHA (MHA) in the particle (class) attention blocks have 8 heads, each with a query dimension of 16 and an expansion factor of 4 for the MLP. We used a dropout (a regularization technique used in neural networks to prevent overfitting) of 0.1 for all P-MAA blocks and no dropout for the class attention block.
5.   Experiment
We perform experiments using the new JETCLASS dataset, a widely adopted benchmark for jet tagging tasks in high-energy physics research. The dataset comprises simulated particle collision events, with each event containing detailed information about the particles produced in the collision, including their momentum and energy. The results of our experiment are summarized in Section 7, where we also compare the performance to previous leading methods.
@@ 6.   Evaluation Metrices
In the context of jet tagging, separating signal jets (class s) from background jets (class b) is crucial. This equation, inspired by the methodologies of the CMS experiment at the LHC, utilizes softmax outputs (Score(s) and Score(b)) to achieve optimal discrimination between the two classes. Background rejection is challenging metric in jet tagging. This is due to significant improvement in background rejection directly translates to a substantial increase in the LHC’s discovery potential. For example, doubling the background rejection can result in a 40% increase in discovery potential, which is comparable to either doubling the dataset size or operating the LHC for twice the duration [^3] The ROC AUC function available in scikit-learn library may be used to calculate the area under the curve by passing in the arguments, average equals to ’macro’ and multi-class equals to ’ovo’..
7.   Results
Evaluation metrics described in Section 6 are used to determine the performance of our proposed model on JETCLASS dataset, with the outcomes compiled in Table 1.This new proposed architecture sets a new benchmark by achieving the highest performance of 86.32%, surpassing all others. ParMAT performs with higher accuracy than Particle transformer.ParMAT significantly outperform the earlier ParT,ParticleNet and P-CNN models on this comprehensive dataset, showing background rejection differences that may exceed an order of magnitude. This significant enhancement in the Particle Multi-Axis Transformer is projected to significantly enhance the discovery capabilities for physics research at the LHC.
@@ 7.   Results
Figure 3: (a) Training Accuracy vs Epochs and (b) Validation Accuracy vs Epochs.
@@
8.   Conclusion
We proposed Multi-axis particle transformer in this paper for jet tagging task on particle collision data from CMS detector. Our model performed well on JETCLASS dataset by achieving a reasonable accuracy for classifying our signals into 10 classes and surpassed many previous ML models proposed for jet tagging including the four baselines models. The AUC of 98.76% reflects the strong discriminatory power of our model and its robustness. It also shows the model’s ability to correctly identify our signals while minimizing the false positives. It also suggest the ability of our model for jet tagging task where precise classification is crucial and false positives can lead to significant drawbacks. Ranger optimizer performed better than Adam for training this model (Complete details about training parameters have been covered in section 5: Experiment). Background rejection matrix shows that our model significantly improved on background noise rejection for both decay modes of top quarks $(t\xrightarrow{}bqq^{\prime}$ $t\xrightarrow{}blv)$ and one decay mode of Higgs$(H\xrightarrow{}gg)$. This improvement resulted from integration of multi axis attention block which uses multi axis attention mechanism to explore relations between different particles and jets by considering multiple axis at a time. Results support the fact that transformers can indeed be a good solution for classification tasks in particle physics where we have large amounts of data generated by detectors. In the future, jet tagging in particle physics experiments is expected to evolve significantly[45]. Researchers can harness particle level information, using individual particle data within jets, including track information and calorimeter shower shapes, to enhance tagging accuracy. Additionally, the global event context can play a crucial role, with algorithms incorporating event-wide data such as missing energy and jet multiplicity to refine tagging strategies. Time series analysis can be employed to distinguish jets in complex events with multiple proton-proton collisions, while calorimeter substructure techniques can be refined to exploit the intricate details of jet formation for improved tagging power.
@@ Acknowledgements
We express our gratitude to the Pattern Recognition Lab (PR Lab) and the Pakistan Institute of Engineering and Applied Sciences (PIEAS) for offering the essential computational facilities and fostering a conducive research atmosphere.
@@
The authors affirm that they have no known conflicts of interest, whether financial or personal, that could have influenced the research presented in this article.